# Maximizing Asynchronicity in Event-based Neural Networks

### v1

@@ Appendix A Architecture details of the asynchronous encoder
This is equivalent to directly using the multi-channel MVHS of the TM in RWKV-6 as the output. The hidden state preserves the aggregated information of the input events, which is precisely what we want for event representation. Moreover, the multi-channel matrix shape matches the multi-channel image-like handcrafted event representations and could convey expressive fine-grained spatial information. Therefore, we think using MVHS as output is a key towards expressive event representation.
Appendix B Dataset preprocess
@@
N-Cars. N-Cars has files of duration 100 ms. Each file has around 4000 events. After splitting the files into patches of size $16\times 16$, the average length of each patch is around 200, which is quite small. Therefore, on this dataset, we split the patches into samples of length 768 with a stride of 512. The first 512 events are inputs, and the last 256 events are used for NRP objectives. Following [12, 5], we use 20% of the training set as the validation split. The samples are saved in the same format as DVS128-Gesture.
Gen1. We first filter out all the hot pixels caused by event camera sensor errors in the dataset. For every 10 ms in the file, the events in a pixel are deleted if they activate more than 40 times (we check the normal files and find this value is no more than 10 regularly). This is similar to limiting the maximum value in the event counts in the synchronous method [14]. We use the official train-val-test split of Gen1 and split the files into patches of size $16\times 16$. Since Gen1 is a large dataset with sufficient events for pretraining the model, we do not use all the events for pretraining in one epoch. Instead, we uniformly sample a slice of length 4096 from each patch. The first 2048 events are inputs, and the last 2048 events are left for NRP objectives. For pretraining, we save the patches in the same format as DVS128-Gesture. For training, all the generated representations are multiplied by a factor of 1/8, converted to their absolute value of data type uint8 to be consistent with the numerical range of the original inputs of RVT (0 to 20). We do this because we find that if we use the original float data type of the representation, there will probably be a numerical NAN error caused by the YOLOX head [11]. This could also help save storage.
Appendix C Implementations
[…]
@@ Appendix C Implementations
In the recognition task on DVS128-Gesture, we use a 14-layer (ResNet-14) as the classifier by removing the last stage of the original ResNet-18 [18] to reduce inference latency (see Table 7). Interestingly, ResNet-14 has a higher accuracy on the validation set with a smaller parameter size and lower latency. We think this is because a smaller model could help overcome overfitting. Despite this, we find that ResNet-18 could also outperform previous work [30]. We use batch size 32, Adam optimizer with learning rate 0.001, mixup augmentation, and train 200 epochs on a single RTX 3090 GPU with a decay of learning rate $\gamma=0.8$ every 20 epochs.
@@
In the experiments on N-Cars, we follow [5, 12] to use a ResNet-34 pretrained on ImageNet as the classifier. Following the train settings of previous work, we use a batch size of 64, Adam optimizer with a constant learning rate of 0.0001, random flip as data augmentation, and train 100 epochs on an RTX 3090 GPU. In the experiments on Gen1, we use the default training setups of RVT[14] for a fair comparison. We use batch size 8, learning rate 2e-4, and train 400,000 steps on a single A800 GPU. The training takes about 2 days.
Appendix D Pretraining objectives
We use event count (EC) and time surface (TS) as the pretraining objectives. Consider events =$\mathcal{E}=\{e_{i}=(t_{i},x_{i},y_{i},p_{i})\}_{i=1}^{T}$ and patch size $P$, EC of a given time window $[t_{s},t_{e}]$ is constructed as a 2 channel image of shape $2\times P\times P$ by
@@ Appendix D Pretraining objectives
100 100 20 0.366 3.16×$\times 10^{-3}$ 0.767 98.1% 94.7%
100 $\rightarrow$ 50 100 20 0.200 2.33×$\times 10^{-3}$ 0.764 95.5% 94.3%
100 $\rightarrow$ 200 100 20 0.807 3.60×$\times 10^{-3}$ 0.767 95.5% 94.3%
100 100 $\rightarrow$ 50 20 0.398 2.39×$\times 10^{-3}$ 0.766 98.1% 94.1%
100 100 $\rightarrow$ 200 20 0.374 3.78×$\times 10^{-3}$ 0.780 96.2% 94.4%
100 100 20 $\rightarrow$ 10 0.389 3.29×$\times 10^{-3}$ 0.364 97.4% 94.2%
100 100 20 $\rightarrow$ 40 0.353 2.86×$\times 10^{-3}$ 1.698 98.7% 94.3%
@@
Appendix E Visualization
We visualize the handcrafted representations for SSL pretraining, the predicted results of the pretraining target, and the learned representations in Figure 5. The results are generated on DVS128-Gesture. The comparison between Figure 5 (a) and (b) demonstrates that during SSL, our model could consistently predict multiple handcrafted representations (EC, TS) of the events with high similarity. Additionally, it shows favorable visualization results in predicting the EC for future time periods. This indicates that our model is capable of predicting short-term motion patterns rather than merely memorizing the history.
@@ Appendix E Visualization
Figure 5: Visualization on DVS128-Gesture. (a) Ground truth of the handcrafted representations. (b) Predicted results of the handcrafted representations. (c) The learned representations.
@@
Appendix F Details of the MVHS and patch-wise encoding to reduce model size
@@
Assume that we keep the size of the learned representation unchanged. For original vector-value output, the output size is $D$, and the model parameter size is around $\mathcal{O}(D^{2})$ (since the majority of the parameters come from the matrix). For MVHS output, the output size is $N\times(D^{\prime}/N)\times(D^{\prime}/N)$ with model size $\mathcal{O}(D^{\prime 2})$. Let $D=N\times(D^{\prime}/N)\times(D^{\prime}/N)$, we have $\mathcal{O}(D^{2})/\mathcal{O}(D^{\prime 2})=D/N=D_{\text{head}}$. In our EVA, this factor is 8 or 16.
@@
For patch-wise encoding, assume that we learn a representation that is proportional to the size of the objective representations. Consider an event camera with resolution $(H,W)$ (assume $H=W$ for simplicity) and patch size $P$, the number of patches #$\#\;patches$ is $H^{2}/P^{2}$. Given the shape of the learned representation as $(N\times D\times D)$, if we use it to predict the objective representation of full size, which is (×$(*\times H\times H)$, then we have $D=H$ and the model size is around $\mathcal{O}(H^{2})$. However, if we use the learned representation to predict a the representations of a patch, which have the shape of (×$(*\times P\times P)$, we have $D=P$, and the model size is around $\mathcal{O}(P^{2})$. Therefore, the model size decreases $\mathcal{O}(H^{2})/\mathcal{O}(P^{2})=\#\;patches$ times. In the experiment on DVS128-Gesture, we have $H=W=128$ and $P=16$. This factor is 64. On Gen1 with $H=240,W=304$ and $P=16$, we have #$\#\;patches=285$.
Appendix G Limitations

### v2

@@ Appendix B Architecture details of the asynchronous encoder
This is equivalent to directly using the multi-channel MVHS of the TM in RWKV-6 as the output. The hidden state preserves the aggregated information of the input events, which is precisely what we want for the event feature. Moreover, the multi-channel matrix shape matches the multi-channel image-like handcrafted event representations and could convey expressive fine-grained spatial information. Therefore, we think using MVHS as output is a key towards expressive event features. The data flow of the EVA framework, from raw events to the targets, is summarized in Figure 6.
Figure 6: Data flow of the framework.
Appendix C Dataset preprocess
@@
N-Cars. N-Cars has files of duration 100 ms. Each file has around 4000 events. After splitting the files into patches of size $16\times 16$, the average length of each patch is around 200, which is quite small. Therefore, on this dataset, we split the patches into samples of length 768 with a stride of 512. The first 512 events are inputs, and the last 256 events are used for NRP objectives. Following (Gehrig et al., 2019; Cannici et al., 2020), we use 20% of the training set as the validation split. The samples are saved in the same format as DVS128-Gesture.
Gen1. We first filter out all the hot pixels caused by event camera sensor errors in the dataset. For every 10 ms in the file, the events in a pixel are deleted if they activate more than 40 times (we check the normal files and find that this value is no more than 10 regularly). This is similar to limiting the maximum value in the event counts in the synchronous method (Gehrig and Scaramuzza, 2023). We use the official train-val-test split of Gen1 and split the files into patches of size $16\times 16$. Since Gen1 is a large dataset with sufficient events for pretraining the model, we do not use all the events for pretraining in one epoch. Instead, we uniformly sample a slice of length 4096 from each patch. The first 2048 events are inputs, and the last 2048 events are left for NRP objectives. For pretraining, we save the patches in the same format as DVS128-Gesture. For training, all the generated features are multiplied by a factor of 1/8, converted to their absolute value of data type uint8 to be consistent with the numerical range of the original inputs of RVT (0 to 20). We do this because we find that if we use the original float data type of the features, there will probably be a numerical NaN error caused by the YOLOX head (Ge et al., 2021). This could also help save storage.
Appendix D Implementations
Three asynchronous encoders are presented in the experiment. In the experiments on DVS128-Gesture, we present an encoder (EVA) with 3 RWKV-6 blocks and 1 MVHS layer with $D=128,D_{\text{head}}=8,N=16,D_{\text{ffn}}=256,D_{\text{lora}}=D_{w}=16$. The shape of the MVHS output is $(16\times 8\times 8)$. To predict the representation of shape $(*\times 16\times 16)$, we use a ResNet head with 1 basic ResNet block of width 64, 1 $\operatorname{ConvTranspose2d}(in=64,out=32,ks=4,stride=2)$ to upsample the $8\times 8$ hidden to $16\times 16$ and 1 basic ResNet block of width 32. We use one $1\times 1$ convolution layer as output. In the experiment on N-Cars, except for the encoder (EVA) on DVS128-Gesture, we also present an encoder with $N=8,D_{\text{head}}=16$ (referred to as EVA ($D_{\text{head}}=16$) in the main context) since the spatial resolution of files in N-Cars is already cropped and small in spatial resolution. Using a shrunk hidden state as feature could indeed help reduce downstream tasks inference latency, but will decrease the accuracy (see Table 2). To keep the same total size of the feature, we use only half the channels in the MVHS layer, that is $N=4,D_{\text{head}}=16$. The shape of the MVHS output is $4\times 16\times 16$. In this case, since there is no need for upsampling, we use a ResNet with 1 basic ResNet block of width 32 as the head. In the experiment on Gen1, we also present a larger encoder with $D=192,D_{\text{head}}=16,N=8,D_{\text{ffn}}=384$ (referred to as EVA-L ($D=192,D_{\text{head}}=16$) in the main context) in the RWKV-6 blocks and $N=6,D_{\text{head}}=16$ in the MVHS layer. The same ResNet head as N-Cars is used for pretraining.
@@ Appendix D Implementations
In the recognition task on DVS128-Gesture, we use a 14-layer (ResNet-14) as the classifier by removing the last stage of the original ResNet-18 (He et al., 2016) to reduce inference latency (see Table 12). Interestingly, ResNet-14 has a higher accuracy on the validation set with a smaller parameter size and lower latency. We think this is because a smaller model could help overcome overfitting. Despite this, we find that ResNet-18 could also outperform previous work (Martin-Turrero et al., 2024). We use batch size 32, Adam optimizer with learning rate 0.001, mixup augmentation, and train 200 epochs on a single RTX 3090 GPU with a decay of learning rate $\gamma=0.8$ every 20 epochs.
@@
In the experiments on N-Cars, we follow (Cannici et al., 2020; Gehrig et al., 2019) to use a ResNet-34 pretrained on ImageNet as the classifier. Following the train settings of previous work, we use a batch size of 64, Adam optimizer with a constant learning rate of 0.0001, random flip as data augmentation, and train 100 epochs on an RTX 3090 GPU. In the experiments on Gen1, we use the default training setups of RVT(Gehrig and Scaramuzza, 2023) for a fair comparison. We use batch size 8, learning rate 2e-4, and train 400,000 steps on a single A800 GPU. The training takes about 2 days.
Appendix E Pretraining objectives
We use event count (EC) and time surface (TS) as the pretraining objectives. Consider events $\mathcal{E}=\{e_{i}=(t_{i},x_{i},y_{i},p_{i})\}_{i=1}^{T}$ and patch size $P$, EC of a given time window $t_{s},t_{e}]$ is constructed as a 2 channel image of shape $2\times P\times P$ by
@@ Appendix E Pretraining objectives
Calculating the loss of the predictions and targets at every individual event would demand prohibitive GPU memory. To mitigate this, we divide the event sequence into chunks, and calculate the loss only at the end of each chunk. For DVS128-Gesture dataset, we use chunk length $T_{\text{chunk}}=512$. For N-Cars and Gen1, we use chunk length $T_{\text{chunk}}=16$ .
@@
[…]
@@
Appendix F Visualization
@@
We visualize the handcrafted representations for SSL pretraining, the predicted results of the pretraining target, and the learned feature in Figure 7. The representations and results are generated on DVS128-Gesture.
Figure 7 (a) provides the visualization of handcrafted representations (targets in SSL), and Figure 7 (b) shows the visualization of the reconstructed representations (predictions in SSL). The comparison between Figure 7 (a) and (b) demonstrates that during SSL, our neural network heads could consistently predict the corresponding handcrafted representations (EC, TS) with high similarity from the learned feature map. This shows that the learned features contain adequate information to reconstruct the handcrafted representations, and further shows that the EVA encoder is able to capture the majority of the spatial information from raw individual events. Additionally, the predicted EC (NRP) shows favorable consistency with its ground truth targets, demonstrating the success of predicting future values from the learned feature map. This indicates that our model is capable of learning short-term motion patterns rather than merely memorizing the inputs.
Figure 7 (c) provides the visualization of the learned feature map. Each figure is one channel of the MVHS layer output, and is a concatenation of multiple patches. Qualitatively, several channels of the learned feature map show patterns similar to the handcrafted representations, but with different windows of temporal accumulation (several channels preserve more past events, while others only maintain recent events). Besides, some channels show different patterns compared with the handcrafted representations. We leave a comprehensive quantitative analysis of the dynamics of the learned feature maps for further study.
Figure 7: Visualization of 4 different gestures on DVS128-Gesture. (a) Ground truth of the handcrafted representations. From left to right: EC, TS, and EC in the future. (b) Neural network head predicted handcrafted representations in the pretraining stage. (c) The learned features. Each figure is one channel of the MVHS layer output.
@@
Appendix G Details of the MVHS and patch-wise encoding to reduce model size
@@
Assume that we keep the size of the learned feature unchanged. For original vector-value output, the output size is $D$, and the model parameter size is around $\mathcal{O}(D^{2})$ (since the majority of the parameters come from the matrix). For MVHS output, the output size is $N\times(D^{\prime}/N)\times(D^{\prime}/N)$ with model size $\mathcal{O}(D^{\prime 2})$. Let $D=N\times(D^{\prime}/N)\times(D^{\prime}/N)$, we have $\mathcal{O}(D^{2})/\mathcal{O}(D^{\prime 2})=D/N=D_{\text{head}}$. In our EVA, this factor is 8 or 16.
@@
For patch-wise encoding, assume that we learn a feature that is proportional to the size of the objective representations. Consider an event camera with resolution $(H,W)$ (assume $H=W$ for simplicity) and patch size $P$, the number of patches #patches\#\;patches is $H^{2}/P^{2}$. Given the shape of the learned feature as $(N\times D\times D)$, if we use it to predict the objective representations of full size, which is $(*\times H\times H)$, then we have $D=H$ and the model size is around $\mathcal{O}(H^{2})$. However, if we use the learned feature to predict the representations of a patch, which has the shape of $(*\times P\times P)$, we have $D=P$, and the model size is around $\mathcal{O}(P^{2})$. Therefore, the model size decreases $\mathcal{O}(H^{2})/\mathcal{O}(P^{2})=\#\;patches$ times. In the experiment on DVS128-Gesture, we have $H=W=128$ and $P=16$. This factor is 64. On Gen1 with $H=240,W=304$ and $P=16$, we have #patches=285\#\;patches=285.
Appendix H Limitations
@@
Appendix I Complexity analysis
Each update of the EVA features requires $\mathcal{O}(M)$ computations, where $M$ is the parameter size of the EVA encoder. Therefore, after a sequential length of T, the total MACs is proportional to the input length $T$ and the model size $M$ as $\mathcal{O}(TM)$. Here we summarize the complexity of other event-representing methods in Table 15. For EST, each time a new event arrives, the representations should be recomputed. The complexity is proportional to the number of events involved in the model size $M$, time window $L$, and number of channels of the EST representation $C$. For MatrixLSTM with temporal bins, the complexity is $\mathcal{O}(LM)$, which serves as a synchronous representation. For the Voxel grid, the complexity is $\mathcal{O}(L+HWC)$, and for ERGO, the complexity is $\mathcal{O}(LC+HWC)$.
Table 15: Complexity and estimated FLOPs of different representations. We consider a $L=1\times 10^{5},H=720,W=1280$ event sequence on the 1Mpx dataset.
Method Asynchronous Complexity per Event Estimated FLOPs per Event
EVA-L $\mathcal{O}(M)$ 1.4 M
ALERT-Tr. (RM) $\mathcal{O}(M)$ 1.2 M
EST $\mathcal{O}(LMC+HWC)$ 1152 M
MatrixLSTM $\mathcal{O}(LMC)$ >> 100 M
Voxel grid $\mathcal{O}(L+HWC)$ 11 M
ERGO-12 $\mathcal{O}(LC+HWC)$ 11 M
Appendix J LLM usage
Large language models (LLMs) are used to improve the writing of the manuscript, including translating languages, correcting grammatical and spelling errors, and improving sentence clarity.